OpenStudy (anonymous):

If f(x) = x2 + 3, find f(5)

OpenStudy (moonlitfate):

All you have to do is plug in a 5 wherever you see an x in your equation and solve it! :) So, in the case of f(x) =2x +3 --> f(5) = 2(5) +3 f(5) = 10+3 = 13, so f(5) = 13. Does that help?

OpenStudy (anonymous):

Yes, thank you!!

OpenStudy (moonlitfate):

You're welcome!